# Equivariant Filter for Tightly Coupled LiDAR-Inertial Odometry*

### v1

@@ Abstract
Pose estimation is a crucial problem in simultaneous localization and mapping (SLAM). However, developing a robust and consistent state estimator remains a significant challenge, as the traditional extended Kalman filter (EKF) struggles to handle the model nonlinearity, especially for inertial measurement unit (IMU) and light detection and ranging (LiDAR). To provide a consistent and efficient solution of pose estimation, we propose Eq-LIO, a robust state estimator for tightly coupled LIO systems based on an equivariant filter (EqF). Compared with the invariant Kalman filter based on the $\operatorname{SE}_{2}(3)$ group structure, the EqF uses the symmetry of the semi-direct product group to couple the system state including IMU bias, navigation state and LiDAR extrinsic calibration state, thereby suppressing linearization error and improving the behavior of the estimator in the event of unexpected state changes. The proposed Eq-LIO owns natural consistency and higher robustness, which is theoretically proven with mathematical derivation and experimentally verified through a series of tests on both public and private datasets.
@@ I Introduction
As a result, LiDAR-inertial odometry (LIO) has been widely used in industry. Among various approaches, filter-based methods are well-suited for low-cost platforms with limited computing resources, as they can efficiently handle large volumes of real-time measurements while maintaining lower computational demands. Among filter-based methods, the extended Kalman filter (EKF) achieves high efficiency, albeit with some loss of accuracy. This trade-off makes it a classic data fusion method, as seen in applications like [4] and others. However, researchers have identified that EKF may yield inconsistent estimates, where the computed covariance deviates from the true covariance [5]. To alleviate the inconsistency problem, Barrau and Bonnabel proposed the invariant extended Kalman filter (IEKF) [6]. IEKF ensures the consistency of the bias-free IMU system by analyzing the system’s group affine properties. Later, Shi et al. [7] applied the IEKF for fusing IMU and LiDAR measurements to reduce estimation error, but the IMU bias will destroy the symmetry of the system. Mahony et al. proposed a modern method known as equivariant filtering (EqF) [8, 9], which naturally enhances consistency and robustness with lower computational overhead by incorporating different system symmetries and fixed linearization points. In light of these advantages, our work fully leverages EqF.
@@ II-B Kalman Filtering for Navigation Applications
Since the introduction of the classic extended Kalman filter (EKF), it has become a widely used tool in filter-based SLAM algorithms. However, the traditional EKF-based LIO algorithm is susceptible to overconfidence issues due to spurious information gain in unobservable directions. To address the inconsistency issues inherent in EKF, Barrau and Bonnabel proposed the invariant extended Kalman filter (IEKF) [6], demonstrating that a consistent filtering algorithm can be developed on group affine systems. This method assumes that the IMU is bias-free, which could not be ignored in practical applications. Despite this limitation, many researchers [15], [16] continue to use this imperfect IEKF framework in the design of SLAM algorithms. They defined the IMU bias on $\mathbb{R}^{3}$, which destroys the group affine property of the system. Recently, SuIn-LIO [17] was proposed, combining an IEKF with an efficient surfel-based map to achieve high-precision LIO. Shi et al.[7] applied the IEKF to the robo-centric and world-centric based LIO and demonstrated that the world-centric method achieves a higher level of accuracy.
Unlike IEKF, which parameterizes the state using Lie groups, Mahony et al. developed an equivariant filtering (EqF) framework [8], [18] that generalizes the state space from Lie groups to homogeneous manifolds. Additionally, EqF does not require constant changes to the linearization point, which naturally ensures system consistency. Subsequently, Fornasier et al. applied EqF to an inertial navigation system [19], which accounts for bias, and demonstrated that various filters, including the EKF, can be incorporated within the EqF framework [20]. Recently, Fornasier et al. [21] introduced an equivariant filter design for vision-assisted INS systems. This new design has been demonstrated to outperform state-of-the-art methods in both robustness and transient behavior.
@@ III-B Group Action, Useful Maps and Notation Explanation
For all $\boldsymbol{X}=(\boldsymbol{C},\boldsymbol{a},\boldsymbol{b})\in\operatorname{SE}_{2}(3)=\operatorname{SO}(3)\ltimes(\mathbb{R}^{3}\oplus\mathbb{R}^{3})$, define $\Gamma$ to extract the rotation part of $\boldsymbol{X}$:
$\Gamma:\operatorname{SE}_{2}(3)\to\operatorname{SO}(3),\quad\Gamma(\boldsymbol{X})=\boldsymbol{C}.$
In this article, the use of bold and non-bold letters for the same character represents two different meanings. Bold letters are generally used to denote matrices, while non-bold letters are typically used to represent elements in groups or manifolds.
@@ IV-A System Definition
where $\boldsymbol{C}_{b}^{w}$, $\boldsymbol{v}_{wb}^{w}$, $\boldsymbol{r}_{wb}^{w}$ represents the attitude, velocity and position of the IMU, $\boldsymbol{b}_{g}$ and $\boldsymbol{b}_{a}$ represent the gyroscope bias and accelerometer bias, respectively, and $\boldsymbol{C}^{b}$ and $\boldsymbol{l}^{b}$ represent the attitude and position of the LiDAR relative to the IMU, $\boldsymbol{g}^{w}$ represents the gravity vector, $\boldsymbol{w}$ and $\boldsymbol{a}$ represent the angular velocity and specific force measured by the IMU.
@@ IV-B Equivariant Symmetry of the System
The symmetry of the system arises from the key semidirect product group $G=\bigl{(}\operatorname{SE}_{2}(3)\ltimes\mathfrak{se}_{2}(3)\bigr{)}\times\operatorname{SE}(3)$. Let $X=(A,a^{\wedge},B)\in G$, $\gamma=(\gamma_{1}^{\wedge},\gamma_{2}^{\wedge},\gamma_{3}^{\wedge},\gamma_{4}^{\wedge})\in L$.
@@
System (4) is equivariant with group action in IV.1 and IV.2, that is,
$f^{0}(\xi)+f_{\psi_{X}(\gamma)}(\xi)\\ =d\phi_{X}\bigl{(}f^{0}(\phi_{X^{-1}}(\xi))+f_{\gamma}(\phi_{X^{-1}}(\xi))\bigr{)}$ (7)
@@ V-A Error Dynamics
For a state $\xi\in\mathcal{M}$, define $e=\phi_{\hat{X}^{-1}}(\xi)\in\mathcal{M}$ to represent the error on the manifold $\mathcal{M}$ [25]. Choose a local coordinate map $\varphi:U\to\mathbb{R}^{9}\oplus\mathbb{R}^{9}\oplus\mathbb{R}^{6}$ at the origin $\xi^{0}$, $\varphi$ can be taken as
$\varphi(e)=(\log^{\vee}e_{1},e_{2}^{\vee},\log^{\vee}e_{3}).$
Let $\varepsilon=\varphi(e)\in\mathbb{R}^{24}$. The linear error $\varepsilon$ satisfies the differential equation
$\dot{\varepsilon}$ $\approx\boldsymbol{F}\varepsilon,$
$\boldsymbol{F}$ $=(d\varphi)_{\xi^{(0)}}\circ\bigl{(}d\phi^{(\xi^{0})}\bigr{)}_{I}\circ\bigl{(}d\Lambda_{\boldsymbol{u}}\bigr{)}_{\xi^{0}}\circ\bigl{(}d\varphi^{-1}\bigr{)}_{\boldsymbol{0}}.$
The closed-form of the matrix $\boldsymbol{F}$ is
$\boldsymbol{F}=\begin{bmatrix}\boldsymbol{F}_{T}&-\boldsymbol{I}_{9\times 9}&\boldsymbol{O}_{9\times 6}\\ \boldsymbol{O}_{9\times 9}&\boldsymbol{F}_{b}&\boldsymbol{O}_{9\times 6}\\ \boldsymbol{O}_{6\times 9}&\boldsymbol{O}_{6\times 9}&\boldsymbol{F}_{K}\end{bmatrix}\in\mathbb{R}^{24\times 24},$ (10)
@@
[…]
$\boldsymbol{Y}=\bigl{(}\hat{\boldsymbol{C}}_{b}^{w}(\boldsymbol{a}-\hat{\boldsymbol{b}}_{a})+(\hat{\boldsymbol{v}}_{wb}^{w})^{\wedge}\hat{\boldsymbol{C}}_{b}^{w}(\boldsymbol{\omega}-\hat{\boldsymbol{b}}_{g})+\boldsymbol{g}^{w}\bigr{)}^{\wedge},$
$\boldsymbol{Z}=\bigl{(}\hat{\boldsymbol{C}}_{b}^{w}(\boldsymbol{\mu}-\hat{\boldsymbol{b}}_{\mu})+(\hat{\boldsymbol{r}}_{wb}^{w})^{\wedge}\hat{\boldsymbol{C}}_{b}^{w}(\boldsymbol{\omega}-\hat{\boldsymbol{b}}_{g})+\hat{\boldsymbol{v}}_{wb}^{w}\bigr{)}^{\wedge},$
$\boldsymbol{W}=\bigl{(}\hat{\boldsymbol{C}}_{b}^{w}(\boldsymbol{\omega}-\hat{\boldsymbol{b}}_{g})\bigr{)}^{\wedge}.$
The letters with hat in them represent the estimated value of that state and $\boldsymbol{I}$ represents the identity matrix.
@@ V-B Measurement Model
At a certain moment, the LiDAR acquires $m$ sampling points $\{\boldsymbol{p}_{j}\}_{j=1}^{m}$. For each point $\boldsymbol{p}_{j}$, We denote point $\boldsymbol{p}_{j}$ transformed to the world coordinate system as $\boldsymbol{p}_{j}^{w}:=\boldsymbol{C}_{b}^{w}(\boldsymbol{C}^{b}\boldsymbol{p}_{j}+\boldsymbol{l}^{b})+\boldsymbol{r}_{wb}^{w}$. Search for the five points closest to $\boldsymbol{p}_{j}^{w}$ and fit a plane, then $\boldsymbol{p}_{j}^{w}$ should be in this plane (Fig. 2). Therefore, the system output $h_{j}:\mathcal{M}\to\mathbb{R}$ can be configured as [14]
@@
$\boldsymbol{H}_{j}=\begin{bmatrix}\boldsymbol{H}_{j,T}&\boldsymbol{O}_{1\times 9}&\boldsymbol{H}_{j,K}\end{bmatrix}\in\mathbb{R}^{1\times 24},$ (11)
@@
$\boldsymbol{H}_{j,T}$ $=\begin{bmatrix}-\boldsymbol{n}_{j}^{\top}{\boldsymbol{p}_{j}^{w}}^{\wedge}&\boldsymbol{O}_{1\times 3}&\boldsymbol{n}_{j}^{\top}\end{bmatrix}\in\mathbb{R}^{1\times 9},$
$\boldsymbol{H}_{j,K}$ $=\begin{bmatrix}-\boldsymbol{n}_{j}^{\top}{\boldsymbol{p}_{j}^{w}}^{\wedge}&\boldsymbol{n}_{j}^{\top}\end{bmatrix}\in\mathbb{R}^{1\times 6}.$
@@ V-C Gravity Constraints on Manifolds $\mathbb{S}^{2}$
We consider gravity to be a vector with constant magnitude and direction on the sphere $\mathbb{S}^{2}$. Since $\mathbb{S}^{2}$ is a two-dimensional manifold, we use a two-dimensional vector in the tangent space $T_{\boldsymbol{g}}\mathbb{S}^{2}\simeq\mathbb{R}^{2}$ as the error state for optimization, which uses the fewest degrees of freedom to estimate gravity.
@@
$B_{\boldsymbol{g}}:\mathbb{R}^{2}\to\mathbb{R}^{3}\quad B_{\boldsymbol{g}}=\begin{bmatrix}1-\frac{x^{2}}{1+z}&-\frac{xy}{1+z}\\[2.84526pt] -\frac{xy}{1+z}&1-\frac{y^{2}}{1+z}\\ -x&-y\end{bmatrix}.$ (12)
$B_{\boldsymbol{g}}$ is an isometry operator, that is, $\left\|B_{\boldsymbol{g}}\boldsymbol{x}\right\|=\left\|\boldsymbol{x}\right\|$ for all $\boldsymbol{x}\in\mathbb{R}^{2}$. Through $B_{\boldsymbol{g}}$, we can convert the vector in the tangent space $T_{\boldsymbol{g}}\mathbb{S}^{2}\simeq\mathbb{R}^{2}$ into a rotation vector, thereby updating the gravity. The pseudoinverse of the matrix $B_{\boldsymbol{g}}$ is $B_{\boldsymbol{g}}^{\top}$.
@@
$\boldsymbol{g}_{k+1}=\exp(B_{\boldsymbol{g}_{k}}\boldsymbol{x})\boldsymbol{g}_{k}\in\mathbb{S}^{2}.$
@@
$\varepsilon_{g}=\arccos(\boldsymbol{g}_{k}\cdot\boldsymbol{g}_{k+1})B_{\boldsymbol{g}_{k}}^{\top}(\boldsymbol{g}_{k}\times\boldsymbol{g}_{k+1}),$ (13)

### v2

@@ Abstract
Pose estimation is a crucial problem in simultaneous localization and mapping (SLAM). However, developing a robust and consistent state estimator remains a significant challenge, as the traditional extended Kalman filter (EKF) struggles to handle the model nonlinearity, especially for inertial measurement unit (IMU) and light detection and ranging (LiDAR). To provide a consistent and efficient solution of pose estimation, we propose Eq-LIO, a robust state estimator for tightly coupled LIO systems based on an equivariant filter (EqF). Compared with the invariant Kalman filter based on the $\operatorname{SE}_{2}(3)$ group structure, the EqF uses the symmetry of the semi-direct product group to couple the system state including IMU bias, navigation state, and LiDAR extrinsic calibration state, thereby suppressing linearization error and improving the behavior of the estimator in the event of unexpected state changes. The proposed Eq-LIO owns natural consistency and higher robustness, which is theoretically proven with mathematical derivation and experimentally verified through a series of tests on both public and private datasets.
@@ I Introduction
As a result, LiDAR-inertial odometry (LIO) has been widely used in industry. Among various approaches, filter-based methods are well-suited for low-cost platforms with limited computing resources, as they can efficiently handle large volumes of real-time measurements while maintaining lower computational demands. Among filter-based methods, the extended Kalman filter (EKF) achieves high efficiency, albeit with some loss of accuracy. This trade-off makes it a classic data fusion method, as seen in applications like [4] and others. However, researchers have identified that EKF may yield inconsistent estimates, where the computed covariance deviates from the true covariance [5]. To alleviate the inconsistency problem, Barrau and Bonnabel proposed the invariant extended Kalman filter (IEKF) by analyzing the system’s group affine properties [6]. Later, Shi et al. [7] applied the imperfect-IEKF for fusing IMU and LiDAR measurements to reduce estimation error, but the IMU bias will destroy the symmetry of the system. Mahony et al. proposed a modern method known as equivariant filtering (EqF) [8, 9], which naturally enhances consistency and robustness with lower computational overhead by incorporating different system symmetries and fixed linearization points. In light of these advantages, our work fully leverages EqF.
@@ II-B Kalman Filtering for Navigation Applications
Since the introduction of the classic extended Kalman filter (EKF), it has become a widely used tool in filter-based SLAM algorithms. However, the traditional EKF-based LIO algorithm is susceptible to overconfidence issues due to spurious information gain in unobservable directions. To address the inconsistency issues inherent in EKF, Barrau and Bonnabel proposed the invariant extended Kalman filter (IEKF) [6], demonstrating that a consistent filtering algorithm can be developed on group affine systems. This method assumes that the IMU is bias-free, which could not be ignored in practical applications. Despite this limitation, many researchers [15], [16] continue to use this imperfect IEKF framework in the design of SLAM algorithms. They defined the IMU bias on $\mathbb{R}^{6}$, which destroys the group affine property of the system. Recently, SuIn-LIO [17] was proposed, combining an IEKF with an efficient surfel-based map to achieve high-precision LIO. Shi et al.[7] applied the IEKF to the robo-centric and world-centric based LIO and demonstrated that the world-centric method achieves a higher level of accuracy.
Unlike IEKF, which parameterizes the state using Lie groups, Mahony et al. developed an equivariant filtering (EqF) framework [8], [18] that generalizes the state space from Lie groups to homogeneous spaces. Additionally, EqF does not require constant changes to the linearization point, which naturally ensures system consistency. Subsequently, Fornasier et al. applied EqF to an inertial navigation system [19], which accounts for bias, and demonstrated that various filters, including the EKF, can be incorporated within the EqF framework [20]. Recently, Fornasier et al. [21] introduced an equivariant filter design for vision-assisted INS systems. This new design has been demonstrated to outperform state-of-the-art methods in both robustness and transient behavior.
@@ III-B Group Action, Useful Maps and Notation Explanation
For all $\boldsymbol{X}=(\boldsymbol{C},\boldsymbol{a},\boldsymbol{b})\in\operatorname{SE}_{2}(3)=\operatorname{SO}(3)\ltimes(\mathbb{R}^{3}\oplus\mathbb{R}^{3})$, define $\Gamma$ to extract the rotation and the position part of $\boldsymbol{X}$:
$\Gamma:\operatorname{SE}_{2}(3)\to\operatorname{SE}(3),\quad\Gamma(\boldsymbol{X})=(\boldsymbol{C},\boldsymbol{b}).$
In this article, the use of bold and non-bold letters for the same character represents two different meanings. Bold letters are generally used to denote matrices, while non-bold letters are typically used to represent elements in groups or manifolds. The proofs of this paper are all available in the supplementary material[22].
@@ IV-A System Definition
where $\boldsymbol{C}_{b}^{w}$, $\boldsymbol{v}_{wb}^{w}$, $\boldsymbol{r}_{wb}^{w}$ represents the attitude, velocity and position of the IMU, $\boldsymbol{b}_{g}$ and $\boldsymbol{b}_{a}$ represent the gyroscope bias and accelerometer bias, respectively, and $\boldsymbol{C}^{b}$ and $\boldsymbol{l}^{b}$ represent the attitude and position of the LiDAR relative to the IMU, $\boldsymbol{g}^{w}$ represents the gravity vector, $\boldsymbol{\omega}$ and $\boldsymbol{a}$ represent the angular velocity and specific force measured by the IMU.
@@ IV-B Equivariant Symmetry of the System
The symmetry of the system arises from the key semidirect product group $G=\bigl{(}\operatorname{SE}_{2}(3)\ltimes\mathfrak{se}_{2}(3)\bigr{)}\times\operatorname{SE}(3)$, which is first introduced by [19]. Let $X=(A,a^{\wedge},B)\in G$, $\gamma=(\gamma_{1}^{\wedge},\gamma_{2}^{\wedge},\gamma_{3}^{\wedge},\gamma_{4}^{\wedge})\in L$.
@@
System (4) is equivariant with respect to group action in IV.1 and IV.2, that is,
$f^{0}(\xi)+f_{\psi_{X}(\gamma)}(\xi)=d\phi_{X}\bigl{(}f^{0}(\phi_{X^{-1}}(\xi))+f_{\gamma}(\phi_{X^{-1}}(\xi))\bigr{)}$
@@ V-A Error Dynamics
Fix the origin state $\xi^{0}=\bigl{(}\boldsymbol{I}_{5\times 5},\boldsymbol{O}_{9\times 9},\boldsymbol{I}_{4\times 4}\bigr{)}$. For current state $\xi\in\mathcal{M}$, define $e=\phi_{\hat{X}^{-1}}(\xi)=\phi_{X\hat{X}^{-1}}(\xi^{0})=\phi_{E}(\xi^{0})\in\mathcal{M}$ to represent the error on $\mathcal{M}$, where $E=X\hat{X}^{-1}$ is right invariant error on $G$ [26]. Choose a local coordinate map $\varphi:U\to\mathbb{R}^{9}\oplus\mathbb{R}^{9}\oplus\mathbb{R}^{6}$ at the origin $\xi^{0}$ , $\varphi$ can be taken as
$\varphi(e)=\log_{G}\circ\bigl{(}\phi^{(\xi^{0})}\bigr{)}^{-1}(e).$
Let $\varepsilon=\varphi(e)\in\mathbb{R}^{24}$. The linear error $\varepsilon$ satisfies
$\dot{\varepsilon}$ $\approx\boldsymbol{F}\varepsilon+\boldsymbol{B}\boldsymbol{n},$
$\boldsymbol{F}$ $=(d\varphi)_{\xi^{(0)}}\circ\bigl{(}d\phi^{(\xi^{0})}\bigr{)}_{I}\circ\bigl{(}d\Lambda_{\boldsymbol{u}}\bigr{)}_{\xi^{0}}\circ\bigl{(}d\varphi^{-1}\bigr{)}_{\boldsymbol{0}},$
where $\boldsymbol{n}$ denotes the Gaussian noise. The state matrix $\boldsymbol{F}$ is
$\boldsymbol{F}=\begin{bmatrix}\boldsymbol{F}_{T}&\boldsymbol{I}_{9\times 9}&\boldsymbol{O}_{9\times 6}\\ \boldsymbol{O}_{9\times 9}&\boldsymbol{F}_{b}&\boldsymbol{O}_{9\times 6}\\ \boldsymbol{F}_{KT}&\boldsymbol{F}_{Kb}&\boldsymbol{F}_{K}\end{bmatrix}\in\mathbb{R}^{24\times 24},$ (9)
@@
$\boldsymbol{F}_{T}$ =[$\ {\boldsymbol{g}^{w}}^{\wedge}&\\ &\boldsymbol{I}_{3\times 3}&\end{bmatrix},\boldsymbol{F}_{b}=\operatorname{ad}_{\Lambda_{1}(\xi^{0},\boldsymbol{u}^{0})}^{\vee}\in\mathbb{R}^{9\times 9},$
$\boldsymbol{F}_{Kb}$ $=\begin{bmatrix}\boldsymbol{I}_{3\times 3}&\boldsymbol{O}_{3\times 3}&\boldsymbol{O}_{3\times 3}\\ \boldsymbol{O}_{3\times 3}&\boldsymbol{O}_{3\times 3}&\boldsymbol{I}_{3\times 3}\end{bmatrix}\in\mathbb{R}^{6\times 9},$
$\boldsymbol{F}_{KT}$ $=\bigl{(}\boldsymbol{I}_{6\times 6}-\operatorname{ad}_{\Gamma(\Lambda_{1}(\xi^{0},\boldsymbol{u}^{0}))}^{\vee}\bigr{)}\begin{bmatrix}\\ &\boldsymbol{I}_{3\times 3}&\end{bmatrix}\in\mathbb{R}^{6\times 9},$
$\boldsymbol{F}_{K}$ $=\operatorname{ad}_{\Gamma(\Lambda(\xi^{0},\boldsymbol{u}^{0}))}^{\vee}\in\mathbb{R}^{6\times 6}.$
The letters with hat in them represent the estimated value of that state and $\boldsymbol{I}$ represents the identity matrix. The derivations of the state matrix $\boldsymbol{F}$ and the input matrix $\boldsymbol{B}$ are privided in the supplementary material [22].
@@ V-B Measurement Model
At a certain moment, the LiDAR acquires $m$ sampling points $\{\boldsymbol{p}_{j}\}_{j=1}^{m}$. Each point $\boldsymbol{p}_{j}$ transformed into the word coordinate system is denoted as $\boldsymbol{p}_{j}^{w}:=(\boldsymbol{K}\Gamma(\boldsymbol{T}))*\boldsymbol{p}_{j}$, where * represents a group action of $\operatorname{SE}(3)$ on $\mathbb{R}^{3}$: $(A,a)*x:=Ax+a$. Search for the five points closest to $\boldsymbol{p}_{j}^{w}$ and fit a plane, then $\boldsymbol{p}_{j}^{w}$ should be in this plane (Fig. 2). Therefore, the system output $h_{j}:\mathcal{M}\to\mathbb{R}$ can be configured as [14]
@@
$\boldsymbol{H}_{j}=\begin{bmatrix}\boldsymbol{O}_{1\times 9}&\boldsymbol{O}_{1\times 9}&\boldsymbol{H}_{j,K}\end{bmatrix}\in\mathbb{R}^{1\times 24},$
@@
$\boldsymbol{H}_{j,K}=\begin{bmatrix}-\boldsymbol{u}_{j}^{\top}\bigl{(}\hat{B}*\boldsymbol{p}_{j}^{w}\bigr{)}^{\wedge}&\boldsymbol{u}_{j}^{\top}\end{bmatrix}\in\mathbb{R}^{1\times 6}.$
@@ V-C Gravity Constraints on Manifolds $\mathbb{S}^{2}$
In order to maintain the minimal parameterization of gravity, we consider gravity to be a vector with constant magnitude and direction on the sphere $\mathbb{S}^{2}$ [27]. Since $\mathbb{S}^{2}$ is a two-dimensional manifold, we use a two-dimensional vector in the tangent space $T_{\boldsymbol{g}}\mathbb{S}^{2}\simeq\mathbb{R}^{2}$ as the error state for optimization, which uses the fewest degrees of freedom to estimate gravity.
@@
$\boldsymbol{B}_{\boldsymbol{g}}:\mathbb{R}^{2}\to\mathbb{R}^{3}\quad\boldsymbol{B}_{\boldsymbol{g}}=\begin{bmatrix}1-\frac{x^{2}}{1+z}&-\frac{xy}{1+z}\\[2.84526pt] -\frac{xy}{1+z}&1-\frac{y^{2}}{1+z}\\ -x&-y\end{bmatrix}.$ (10)
$B_{\boldsymbol{g}}$ is an isometry operator, that is, $\left\|\boldsymbol{B}_{\boldsymbol{g}}\boldsymbol{x}\right\|=\left\|\boldsymbol{x}\right\|$ for all $\boldsymbol{x}\in\mathbb{R}^{2}$. Through $\boldsymbol{B}_{\boldsymbol{g}}$, we can convert the vector in the tangent space $T_{\boldsymbol{g}}\mathbb{S}^{2}\simeq\mathbb{R}^{2}$ into a rotation vector, thereby updating the gravity. The pseudoinverse of the matrix $\boldsymbol{B}_{\boldsymbol{g}}$ is $\boldsymbol{B}_{\boldsymbol{g}}^{\top}$.
@@
$\boldsymbol{g}_{k+1}=\exp(\boldsymbol{B}_{\boldsymbol{g}_{k}}\boldsymbol{x})\boldsymbol{g}_{k}\in\mathbb{S}^{2}.$
@@
$\varepsilon_{g}=\arccos(\boldsymbol{g}_{k}\cdot\boldsymbol{g}_{k+1})\boldsymbol{B}_{\boldsymbol{g}_{k}}^{\top}(\boldsymbol{g}_{k}\times\boldsymbol{g}_{k+1}),$ (11)